OpenStudy (anonymous):

between what two consecutive integers does the square root if 14 lie

OpenStudy (anonymous):

62 lies in between what two integer

OpenStudy (freckles):

62?

sammixboo (sammixboo):

First, what is the square root of 14?

OpenStudy (freckles):

do you mean sqrt(62)?

OpenStudy (freckles):

if we have an integer n, then n is between n-1 and n+1

OpenStudy (freckles):

where n-1 and n+1 are also integers

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\) An exmaple of a perfect square would be something like: 64 because it multiplies 8*8. Same number. @sasharadford578 mind clearing up your question a little? Thank you=)

OpenStudy (freckles):

Anyways for the first question, you could think what two squares is 14 in between? <--as a first step

OpenStudy (freckles):

1^2= 2^2= 3^2= 4^2= 5^2=

OpenStudy (triciaal):

do you have any questions about the post by freckles? I would use the same approach